Using the Mass-Luminosity relationship, estimate the luminosities for each of the following stars:

  1. 25 MSun star
  2. 2.5 MSun star
  3. 0.25 MSun star
Answer #1

luminosity is proportional to 3.5 times of mass so if mass of main sequence star is 2 times the mass of sun then luminosity will be 11.31 of sun.luminosity of sun is 3.83*10^26.
(a)if mass is 25 M sun then luminosity is 78125*luminosity of sun, 78125*3.846*10^26 = 3.0047e31
(b)if mass is 2.5 M sun then luminosity is 24.70*luminosity of sun, 24.7*3.846*10^26 = 9.5*e27
(c)if mass is 0.25 M sun then luminosity is 0.0078125*luminosity of sun, 0.0078125*3.846*10^26 = 3e24